# SLTrain: a sparse plus low-rank approach for parameter and memory efficient pretraining

### v1

@@ 6 Concluding Remarks
In this paper, we propose SLTrain for achieving memory and parameter efficient pretraining of LLMs. SLTrain combines two complementary parsimonious structures, low-rank and sparsity, for learning models with high representation capacity. While low rank is modeled via the product $BA$ of $r$-dimensional matrices, the support of the sparse factor $S$ is obtained by uniformly random sampling over indices. The matrices $B,A,S$ are learned for different layers via backpropagation. Empirically, we achieve state-of-the-art memory reduction during pretraining while maintaining competitive performance. Although we show results on the LLaMA language models (with varying size from 60M to 7B parameters), we believe SLTrain could also help to improve memory efficiency for vision foundation models and multi-modal foundation models, such as diffusion models [37] and CLIP [38]. The significant improvements shown by SLTrain also motivates future works to understand the theoretical guarantees of training with both low-rank and sparse factors, such as convergence and loss landscape. We hope this work initiates exploration on combination of other parsimonious structures for pretraining such as Kronecker product or structured sparsity (e.g., block-diagonal, group-lasso).
Appendix A Additional illustration for low-rank and residual factors
In the main text, we only show a single layer’s singular values along with the attention output weight matrix visualization before and after removing low-rank approximation. Here, we include the visualizations for other weight matrices and other layers for the LLaMA 60M and 130M models.
@@ Appendix A Additional illustration for low-rank and residual factors
Appendix B Singular value distribution of learned weights
@@
Appendix C Details of memory estimation
@@
Appendix D Fine-tuning LLMs
For fine-tuning results, we tune three hyperparmaters of SLTrain, i.e., sparsity level $\delta$, balancing parameter $\alpha$ and stepsize $\eta$. Specifically, we tune $\delta$ in the range of $[0.005,0.001]$ and $\alpha$ in $[32,64,128]$, $\eta$ in [1e-5, 2e-5, 3e-5, 4e-5, 5e-5]. The tuned hyperparameters are in Table 10.

### v2

@@ 6 Concluding Remarks
In this paper, we propose SLTrain for achieving both memory and parameter efficient pretraining of LLMs. SLTrain combines two complementary parsimonious structures, low-rank and sparsity, for learning models with high representation capacity. While low rank is modeled via the product $BA$ of $r$-dimensional matrices, the support of the sparse factor $S$ is obtained by uniformly random sampling over indices. The matrices $B,A,S$ are learned for different layers via backpropagation. Empirically, we achieve state-of-the-art memory reduction during pretraining while maintaining competitive performance. Although we show results on the LLaMA language models (with varying size from 60M to 7B parameters), we believe SLTrain could also help to improve memory efficiency for vision foundation models and multi-modal foundation models, such as diffusion models [37] and CLIP [38]. The significant improvements shown by SLTrain also motivates future works to understand the theoretical guarantees of training with both low-rank and sparse factors, such as convergence and loss landscape. We hope this work initiates exploration on combination of other parsimonious structures for pretraining such as Kronecker product or structured sparsity (e.g., block-diagonal, group-lasso).
Acknowledgments
M. Hong and J. Li are supported partially by NSF under the grants EPCN-2311007 and CCF-1910385, and an Amazon Research Award.
Appendix A Proofs
Proof of Proposition 1.
For each row, the probability that at least one entry is non-zero is $1-(1-p)^{n}$ due to the independence of the non-zero entry. Due to the independence of uniform selection, the probability that all the rows have at least one entry is $(1-(1-p)^{n})^{n}$. By choosing $p=2\log n/n$, the probability can be simplified to $(1-e^{-2\log n})^{n}=1-O(1/n)$. Similarly, we can apply the same argument for the columns. Taking the union bound shows that with probability at least $1-O(1/n)$, $S$ has at least one entry for each row and each column respectively. Further, because the non-zero entries of $S$ are sampled from continuous distribution, the set of matrices such that $S$ becomes low-rank has measure zero. Then taking union bound gives with probability at least $1-O(1/n)$, $S$ is full rank.
The remaining part is to prove that $BA+S$ is full rank. Because $B,A$ are sampled from a continuous distribution space, the set of such matrices that augments $S$ to form a degenerate matrix has measure zero. Thus, the proof is complete. ∎
Appendix B Additional illustration for low-rank and residual factors
In the main text, we present singular values and visualizations of the attention output weight matrix for a single layer, both before and after applying the low-rank approximation. Here, we extend this analysis by providing visualizations for additional weight matrices across multiple layers for the LLaMA 60M and 130M models. The figures reveal a consistent pattern, showing a low-rank plus (uniformly) small magnitude structure in the weight matrices of other layers as well.
@@ Appendix B Additional illustration for low-rank and residual factors
Appendix C Illustration of low-rank and residual factors for Pythia
Here we repeat the analysis of singular spectrum for pretrained Pythia 70M, downloaded from Hugging Face. Specifically, we set the rank $r=128$ and extract the best rank $r$ approximation of the learned weight matrices. The results are shown in Figure 9, where we observe that the residual after removing the best low-rank approximation vary smoothly and has small magnitude.
Figure 9: Illustration of the last layer of pretrained full-rank Pythia 70M (deduped) downloaded from Hugging Face.
Appendix D Singular value distribution of learned weights
@@
Appendix E Memory and runtime of SLTrain linear layer
Here we perform an additional experiment by comparing the actual maximum memory consumption for the proposed SLTrain linear layer (Algorithm 1, $(BA+S)x$) and standard full-rank linear layer ($Wx$) and low-rank linear layer ($BAx$) in a feedforward neural network. We include the results for both forward and backward pass where we vary the number of layers in Figure 112. Specifically, we set the input, hidden and output size to be $2048$ and $r=128$ with $\delta=0.03$. From the figure, we observe that as the number of layers increases, the reduction in memory of SLTrain becomes more evident compared to full-rank model. On the other hand, the memory overhead of SLTrain compared to low-rank model is only marginal. In terms of computational cost, we see compared to full-rank model, SLTrain only requires slight computational overhead, which is due to the scatter-adding operation.
Figure 12: Comparison of memory and runtime with feedforward neural network with full-rank linear layer (FFN), low-rank linear layer ($BA$) and SLTrain linear layer ($BA+S$). We observe the memory savings of SLTrain becomes more evident when number of layers increases. Compared to Low-Rank, the memory of SLTrain is only marginally higher.
Appendix F Details of memory estimation
@@
Appendix G Fine-tuning LLMs
In addition to pretraining, the sparse plus low-rank factor may also be used for fine-tuning LLMs as: $W=W_{0}+BA+S$, where $W_{0}$ is a given pretrained model weight, $B$ and $A$ are low-rank factors, and $S$ is the sparse factor. As in Section 3.2, we learn the fine-tuned weights $B,A,$ and $S$ and term this as SLTrain-FT. The fine-tuning experiments are conducted for RoBERTa base model (with 125M parameters) on GLUE benchmarks. We use $r=8$ for all methods and tune the hyperparameters of SLTrain-FT as follows. We tune three hyperparmaters of SLTrain, i.e., sparsity level $\delta$, balancing parameter $\alpha$ and stepsize $\eta$. Specifically, we tune $\delta$ in the range of $[0.005,0.001]$ and $\alpha$ in $[32,64,128]$, $\eta$ in [1e-5, 2e-5, 3e-5, 4e-5, 5e-5]. The tuned hyperparameters are in Table 11.
The results on benchmark datasets are shown in Table 12. We observe that SLTrain-FT performs competitively as the baselines. While there is no specific memory advantage of SLTrain-FT when a general full-rank $W_{0}$ is given, memory reductions can be obtained when $W_{0}$ is learned via SLTrain.
@@ Appendix G Fine-tuning LLMs
CoLA STS-B MRPC RTE SST-2 MNLI QNLI QQP Avg
Full-rank FT 62.24 90.92 91.30 79.42 94.57 87.18 92.33 92.28 86.28
LoRA (rank=8) 61.83 90.80 91.90 79.06 93.46 86.94 92.25 91.22 85.93
Appendix H Experiment Configurations
We provide the source code for reproducing the experimental results reported in the paper. We also summarize some specific configurations that enhance reproducibility.
  • Datasets. The C4 dataset is publicly available on Huggingface and can be loaded using datasets package https://github.com/huggingface/datasets.
  • Random seed. For all the main experiments for pretraining, we use random seed 42. For fine-tuning experiments, we use random seed 1234.
  • GPU and runtime. All experiments are conducted with multiple runs on NVIDIA A100-SXM4-40GB/80GB GPUs. The training time vary from 2 hrs (LLaMA 60M) to 5 days (LLaMA 7B) depending on the model size. For fine-tuning, the runtime is roughly the same as LoRA, which is reported on https://github.com/microsoft/LoRA/tree/main/examples/NLU/examples/text-classification.
Appendix I Broader Impact
The paper proposes a simple strategy that achieves both parameter and memory efficiency of training LLMs. We believe this work would produce positive environmental impact by reducing the energy consumption as well as carbon footprint during pretraining large foundation models.